OpenStudy (anonymous):

Solve the system of equations and check. 3x - 4y=-15 4x + 3y=5

OpenStudy (anonymous):

I got x=-1

OpenStudy (anonymous):

is this right

OpenStudy (texaschic101):

3x - 4y = -15 --->(-4) 4x + 3y = 5 --->(3) -------------- -12x + 16y = 60 (result of multiplying by -4) 12x + 9y = 15 (result of multiplying by 3) --------------add 25y = 75 -- divide by 25 y = 3 now sub 3 in for y in either equation to find x 3x - 4y = -15 3x - 4(3) = -15 3x - 12 = -15 3x = -15 + 12 3x = -3 x = -1 check.. 4x + 3y = 5 4(-1) + 3(3) = 5 -4 + 9 = 5 5 = 5 (correct) x = -1 and y = 5 any questions ?

OpenStudy (texaschic101):

this is the elimination way....I can do it the substitution way if you need it done that way